OpenStudy (anonymous):

ahmad bought some fish from a farm.during the first week 1/8 of them died while in next two weeks 3/5 of the remaining died.what fraction of the original fish in last two weeks?

OpenStudy (saifoo.khan):

add both!

OpenStudy (anonymous):

Are you looking for the fraction of the remaining fishes or of the dead fishes?

OpenStudy (anonymous):

yes for fraction

OpenStudy (anonymous):

multiply both.

OpenStudy (anonymous):

I got 7/20 for the remaining... Though I'm not sure if it's correct.

OpenStudy (saifoo.khan):

\[\Huge \frac{3}{40}\]

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

its 3/40, just multiply.

OpenStudy (akshay_budhkar):

3/40 is right no????????/

OpenStudy (anonymous):

it is!

OpenStudy (anonymous):

RIGHT ANSWER IS 1/2 NOT 3/40

OpenStudy (akshay_budhkar):

i will stick to 3/40.. 1/2???

OpenStudy (anonymous):

i guess we didnt completely answer the question. in our defense though, didnt tell us what you needed. "What fraction of the original fish in the last two weeks" What fraction of the original fish what? lived? died? gotten eaten by polar bears?

OpenStudy (anonymous):

gotten eaten....got*eaten.

OpenStudy (akshay_budhkar):

lol

OpenStudy (akshay_budhkar):

iman waheed ????

OpenStudy (anonymous):

WHAT U ASKED I DON'T UNDERSTAND

OpenStudy (saifoo.khan):

LOL.

OpenStudy (akshay_budhkar):

"What fraction of the original fish in the last two weeks" What fraction of the original fish what? lived? died? got eaten by polar bears?

OpenStudy (akshay_budhkar):

look at ur question dude

OpenStudy (akshay_budhkar):

joe its d 2 of us only! lol

OpenStudy (anonymous):

it happens lol

OpenStudy (saifoo.khan):

who 2? wat two?

OpenStudy (akshay_budhkar):

u came now saifoo.. but wheres iman?

OpenStudy (saifoo.khan):

Lol, she is busy eating i guess.

OpenStudy (saifoo.khan):

eating fishes i guess.

OpenStudy (anonymous):

i think we did something wrong >.>

OpenStudy (akshay_budhkar):

what??????

OpenStudy (saifoo.khan):

or Joe's watermelon!

OpenStudy (akshay_budhkar):

she didnt specify what she wanted..

OpenStudy (anonymous):

lol, if 1/8 died, that means 7/8 lived right?

OpenStudy (akshay_budhkar):

ya everytime i see joe i ffeel hungry.. lol

OpenStudy (akshay_budhkar):

19/40?

OpenStudy (saifoo.khan):

actually there is are mistakes coz iman is busy eating the fishes.. their amount keeps in changing.

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

then after that, 3/5 of the remaining died, so 2/5 lived, so we get: \[\frac{7}{8}*\frac{2}{5} = \frac{7}{20}\] 7/20 lived, so 13/20 died.

OpenStudy (akshay_budhkar):

i am getting 19/40...:(

OpenStudy (akshay_budhkar):

joe dude its 3/5 th of remaining so 2/5th of 7/8th lived!!!

OpenStudy (akshay_budhkar):

u cant add 2/5???

OpenStudy (anonymous):

I got the same answer as joemath.

OpenStudy (anonymous):

right, its (2/5)*(7/8) (what askhay said)

OpenStudy (akshay_budhkar):

so d ans is 7/20

OpenStudy (akshay_budhkar):

which is what u got.. sorry

OpenStudy (akshay_budhkar):

cwd yes i think we r right

OpenStudy (anonymous):

Joe math how u got answer plz let me explain

OpenStudy (anonymous):

With these types of problems, its important to keep track of what the numbers we are getting represent. That is where I messed up earlier. So lets start from the beginning. In the first week, 1/8 of the fish died. That means that: \[1-\frac{1}{8} = \frac{7}{8}\]lived. Now, in the next two weeks, 3/5 of the fish left also died. That means: \[1-\frac{3}{5} = \frac{2}{5}\]of the 7/8 left alive from the first week survied the next two weeks. This gives us: \[\frac{7}{8}*\frac{2}{5} = \frac{14}{40} = \frac{7}{20}\] of the fish lived after the last 2 weeks. This also means that: \[1-\frac{7}{20} =\frac{13}{20}\]died after the last two weeks. You dont really specify which answer you are looking for in your question, so it could be any one of those 2 answers.